# The SRG/eROSITA All-Sky Survey: X-ray beacons at late cosmic dawn

### v1

@@ 4.3.1 Confirming the X-ray detection with Chandra ACIS-S
Figure 11: All images are centered on the optical position of the new quasar J202040-621509 and have sizes 100’ x100”. Left panel: Broadband (0.5-7 keV) Chandra image. Chandra sources detected with 30” (black, dashed circle) of the quasar coordinates with wavdetect are marked in red. The centroid of 1eRASS J202039.8-621525 catalogue detection is marked by a green triangle. The contours were obtained from the smoothed eRASS:4 0.5 - 7 keV image. Central panel: DES DR2 z-band image.Circular colored markers denote counterpart candidates for 1eRASS J202039.8-621525 from the DESI Legacy Survey catalogue. They are colour-coded according to their relative probability p_i of being the best counterpart. The DESI Legacy Survey source with the highest p_i is spatially co-incident with the quasar coordinates. Right panel: CatWISE 2020 W1 image. Among the Chandra and eROSITA X-ray detections, only the quasar is clearly detected.
@@
We extracted fluxes for the three Chandra sources using the srcflux routine of the ciao software package. For all three sources we assumed a fiducial powerlaw model zpowerlw with Galactic absorption and fixed photon index ($\Gamma=1.9$). We report the broad-band photometry for these three detections in Table 4. ch1 is the most significant detection among the three sources in the viscinity of the quasar. We note that no counts are registered in the ultrasoft band (0.2 - 0.5 keV) and the soft band (0.5-1 keV) for ch1. For this source we measure $5.85\pm 2.45$ net counts in the medium band (1 - 2 keV) and $4\pm 2$ counts in the hard band (2-7 keV). This hard photometry is consistent with a significant level of obscuration, as no counts can be measured in the restframe 1.34 - 6.7 keV energy of the quasar SED.
In summary we have presented evidence that supports that 1eRASS J202039.8-621525 is associated with the quasar emission, as (A) the Chandra sources ch2 and ch3 are marginally detected and do not overlap with the eRASS contours. (B) These sources are also not significantly detected in DES and CatWISE2020. (C) Among 151 potential DESI Legacy Survey candidates within 1’ of 1eRASS J202039.8-621525, the candidate with the highest relative probability p_i of being associated to the X-ray source is matched at the sub-arcsecond level with quasar coordinates. (D) No other Chandra source is detected to the south-west of the quasar, close to eROSITA events. While we cannot exclude the possibility of a transient phenomenon in a foreground galaxy, we conclude that it is likely that the bright eRASS1 source is associated with the quasar.
Source
@@
$F_{0.5-7.0\,\mathrm{keV}}$
[$10^{-15}\mathrm{erg\,cm^{-2}s^{-1}}$]
ch1 0.30 3.06 $5.83^{3.70}_{-2.63}$
ch2 21.57 2.00 $2.86^{2.98}_{-1.82}$
ch3 29.96 2.40 $3.80^{3.20}_{-2.09}$
Table 4: Chandra ACIS-S photometry for three sources detected within 30” of the quasar. [1]The separation in arcsec between the optical position and the Chandra source. [2] the srcflux-based significance of the detection based on source and background counts. [3] The flux in observed 0.5-7 keV.
@@ 4.3.2 Long-term X-ray variability
The eROSITA image presented in Fig. 11 displays two distinct lobes, which extend to the south and south-west of the coordinates of the quasar. One of these lobes coincides astrometrically with the quasar. In the Chandra image, three sources were detected with the wavdetect algorithm. One of these sources is matched at sub-arcsecond level to the quasar’s coordinates. The two other sources are detected at a lower significance level to the north and the north-west of the quasar. Their location does not match the eRASS contours and they are associated with DES sources within 1”. Assuming that the eRASS1 photometry is solely associated with the quasar, the eRASS and Chandra photometry presented respectively in Section 4.1 and 4.3.1, imply a decrease in the measured 0.5 - 7 keV flux of the source of about an order of magnitude from $F_{0.5-7\rm keV}=7.82^{+11.06}_{-4.47}\times 10^{-14}\mathrm{erg\,s^{-1}cm^{-2}}$ to $F_{0.5-7\rm keV}=5.83^{+3.70}_{-2.63}\times 10^{-15}\mathrm{erg\,s^{-1}cm^{-2}}$. We note that the flux decrease is only significant at the $\sim 2\sigma$ level due to the low number of net source counts in eRASS1. The two observations are separated by 3.5 years in the observed frame corresponding to 6 months in the quasar restframe.
Fig. 12 shows the flux-decrease in the broad 0.5-7 keV band. The BXA flux posteriors for the eRASS1 and Chandra , their median and 15.9th and 84.1th percentiles are shown. We have further derived $L_{2-10\,\mathrm{keV}}$ and $\alpha_{OX}$ from the this Chandra flux posteriors. They are respectively marked in the left and right panel of Fig. 8.
@@ 4.3.3 Fast outflows
Here $f(v)$ denotes to the continuum-normalized flux, $C$ is a constant equal to 1 for velocities $>2000\,\mathrm{km\,s^{-1}}$ with respect to the trough edge and 0 elsewhere. The measurement of the continuum is challenging due to the numerous BALs. We have approximated the continuum by using the principal component decomposition of $z\sim 3$ SDSS quasar continua presented by Pâris et al. (2011). The spectrum is strongly attenuated toward longer wavelengths, which is not accounted for by the PCA model. We therefore fitted an additional powerlaw in the restframe spectral regions 1300-1350 $\AA$ and 1400-1600 $\AA$ to the spectrum normalized by the PCA-reprojected model. Finally, the spectrum of the source was normalised by the combined PCA and powerlaw models. We account for the flux uncertainties by bootstrapping from the spectrum RMS. From the normalized spectrum we measure $BI_{\mathrm{CIV}}$ = $3394\pm 197\,\mathrm{km\,s^{-1}}$ and $BI_{\mathrm{SiIV}}$ = $2935\pm 110\,\mathrm{km\,s^{-1}}$. This BAL strengths confirm the classification of J202040-621509 as BAL quasar.
The large BALs are detached with respect to the quasar redshift, i.e. the troughs of all species are shifted shortwards of the expected corresponding quasar emission lines. We measure detached and terminal velocities from the start and the end of the absorption troughs. From the SiIV BAL we measure outflow velocities $v_{\mathrm{SiIV}}=0.002c-0.020c$ and $v_{\mathrm{CIV}}=0.003c-0.025c$.
@@ 4.4 Radio properties
J202040--621509 is clearly detected at radio frequencies, at 0.887 and 1.37 GHz in VAST and at 0.943 GHz in the EMU Pilot survey (Norris et al. 2011). We retrieved the corresponding radio images from CASDA and we performed a single Gaussian fit using the task IMFIT within the CASA software to quantify the flux densities (see Table 6). The images are shown in Fig. 15 (top two rows) in Appendix C. Several VAST images are available, corresponding to different observing dates. We extracted the flux density from each of them. By assuming a single power-law for the continuum radio emission, we obtained a radio spectral index ($\alpha_{r}$) of 0.07$\pm$0.15 (see Fig. 14) that makes J202040--621509 FSRQ, similar to J020916--562650. The VAST measurements are consistent with each other, indicating that the source does not appear to be variable on time scales of a few months (observed frame). Then we computed the values of radio-loudness (R), by following the same procedure as for J020916--562650. We obtained a mean R of 6.5$\pm$0.6, which formally classifies the source as radio-quiet, since the standard separation between radio-loud and radio-quiet objects is set at R=10 (Urry & Padovani 1995). However, this does not rule out the presence of a relativistic jet (e.g., Sbarrato et al. 2021).
@@
4.4.1 Obscuring fast accretion disc winds at high redshift and other variability scenarios
@@
AGN shows substantial variability in flux and spectral shape at all frequencies (for a review, see Lira 2021). This variability arises from accretion physics (e.g. Nandra et al. 1997), broad line region kinematics (e.g. Cherepashchuk & Lyutyi 1973; Boksenberg et al. 1978; Blandford & McKee 1982; Peterson 1993) and jets (Albert et al. 2007, e.g.). In X-rays, AGN has been observed to be variable on timescales of hours to days (e.g. Winkler & White 1975; Marshall et al. 1981; McHardy et al. 2004; Ponti et al. 2012). Little redshift evolution in the X-ray variabity of AGN, as traced by the variance-frequency diagram, has been observed from $z=0$ to $z=3$ (Paolillo et al. 2023). Interpreted as reverberation effect, i.e. reprocessing of accretion disk photons in the electron corona (Haardt & Maraschi 1991), rapid X-ray variability can be used to study the structure of the innermost region of the AGN, at couple gravitational radii away from the black hole (e.g. McHardy 2013). Coronal variability usually shows maximum amplitudes of a factor $\sim 2$ (e.g., Gibson & Brandt 2012). However, more extreme variability, unrelated to coronal fluctuations, has been observed on longer timescales (e.g. Timlin et al. 2020). For instance, (LaMassa et al. 2015) reported the first discovery of a changing look quasar in which the AGN flux dropped by a factor of 6 over timescales of years while not showing absorption signatures. As presented in Section 4.3.3, the UV spectrum of J202040-621509 in contrast is that of a typical BAL quasar. The X-ray variability in quasars containing BALs in their restframe UV spectra is thought to arise from absorption effects. More precisely variations in the column density of intervening, obscuring clouds are expected to arise from large-scale outflows (e.g. Brandt et al. 2000; Green et al. 2001; Gallagher et al. 2002, 2006; Gibson et al. 2009; Saez et al. 2012). An often invoked physical model for the observed extreme outflow velocities, is that of line-driven winds launched from the accretion disk, as high-accretion rates drive the UV radiation pressure of the disk up. High-density regions, located at the inner edge of this flow can shield the coronal X-ray emission through photo-electric absorption (NH$>10^{22}\,\rm cm^{-2}$ Murray et al. 1995).
@@
A possible scenario is that the observed X-ray dimming of J202040-621509 over one order of magnitude in X-ray flux may be in part explained by such obscuring winds. While too few counts have been collected in the Chandra observation to extract and fit an X-ray spectrum, we only collected counts at energies $>6.7\,\rm keV$ in the quasar restframe. This hard photometry is indicative of intrinsic X-ray absorption, affecting lower-energetic photons.
@@
In the early universe, X-ray variability has previously been reported in two objects (Nanni et al. 2018; Vito et al. 2022). Nanni et al. (2018) reported a hardening of the spectrum and a flux decrease of a factor 2.5 in the quasar SDSS J1030+0524 between two XMM-Newton observations spaced by 14 years. They proposed intrinsic variation of accretion rate as potential driver of the observed variability. Vito et al. (2022) presented the surprising non-detection in a 100ks XMM-Newton observations of the quasar CFHQS J164121+375520, which had been detected 115 quasar-restframe days before with Chandra, implying a flux-dimming of factor $>7$. Intervening obscuring material on sub-pc scales such as disk-shape variability in super-Eddington sources is discussed by the authors. Vito et al. (2022) also list a possible model by Giustini & Proga (2019), in which rapidly accreting black holes with $\dot{m}\geq 0.25$, generate the favourable conditions for their optically-thick, thin disk emission to dominate the output of a cooled corona. These disks are expected to launch strong line-driven winds, imprinting BAL features in the spectrum of the quasars.
@@
While the latter model yields a promising scenario for the phenomonology observed in J202040-621509, we cannot exclude any of the above mentioned sources of variability as the counting statistics of our X-ray observations are poor and we are missing simultaneous UV photometry to simultaneously track the corona and disk emission. A deeper, simultaneous X-ray and UV monitoring campaign will be required to shed light on this seeminlgy extreme X-ray dimming event.
@@ 5 Conclusions
In this work, we have presented a pilot survey, designed to uncover new X-ray luminous quasars at $z>5.6$ combining eROSITA data with optical and IR photometry from DES, VHS and CatWISE2020. We have discovered 5 quasars out of 9 sources, that were selected for spectroscopic confirmation with LDSS3-C. Two of these sources are securely detected in eRASS. These source are also radio-detected (and radio-loud). The reionisation-era quasars uncovered by our pilot study shed a light on powerful X-ray emission mechanisms in the early universe and gives us access to demographic diagnostics for the $z>5.6$ quasar population.
J020916-562650 ($z=5.6$) shows X-ray and radio properties of a blazar, making it the second most distant AGN of this type known to date. From this detection and using simple geometric arguments, we infer a space density of similarly luminous, jetted quasars of $6.89^{15.86}_{-5.70}\,\mathrm{Gpc^{-3}}$. This value is consistent with the space density of all quasars, jetted or not, as measured by the most recent quasar luminosity functions at this redshift and brightness. Our measurement implies two potential scenarios: either all quasars discovered so far at $z\sim 6$ and M1450¡-25 are jetted (and most of these jets have not been detected yet) or a substantial fraction of jetted quasars in this redshift-magnitude regime have eluded quasar searches so far, perhaps due to obscuration. Severe systematics that can affect our measurement of quasar space density are discussed in this paper.
J202040-621509, ($z=5.7$) is a BAL quasar. Through four consecutive eRASS observations and an additional Chandra pointing, we provide substantial evidence for an extreme X-ray dimming event (a factor $\geq 13$ in X-ray flux). Potential dimming scenarios are nuclear obscuration or disk dominated quasar SEDs in strongly outflowing quasars with rapidly accreting black holes. We conclude, however, that increased X-ray counting statistics and simultaneous UV monitoring will be required to draw clearer conclusions on this dimming event.

### v2

@@ 4.3.1 Confirming the X-ray detection with Chandra ACIS-S
Figure 11: All images are centred on the optical position of the new quasar J202040-621509 and have sizes 100” x100”. Left panel: Broadband (0.5-7 keV) Chandra image. Chandra sources detected with 30” (black, dashed circle) of the quasar coordinates with wavdetect are marked in red. The centroid of 1eRASS J202039.8-621525 catalogue detection is marked by a green triangle. The contours were obtained from the smoothed eRASS:4 0.5 - 7 keV image. Central panel: DES DR2 z-band image. Circular coloured markers denote counterpart candidates for 1eRASS J202039.8-621525 from the DESI Legacy Survey catalogue. They are colour-coded according to their relative probability p_i of being the best counterpart. The DESI Legacy Survey source with the highest p_i is spatially co-incident with the quasar coordinates. Right panel: CatWISE 2020 W1 image. Among the Chandra and eROSITA X-ray detections, only the quasar is unambiguously detected.
@@
We extracted fluxes for the three Chandra sources using the srcflux routine of the ciao software package. For all three sources, we assumed a fiducial power-law model zpowerlw with Galactic absorption and fixed photon index ($\Gamma=1.9$). We report the broad-band photometry for these three detections in Table 10. ch1 is the most significant detection among the three sources near the quasar. We note that no counts are registered in the ultrasoft band (0.2 - 0.5 keV) and the soft band (0.5-1 keV) for ch1. For this source, we measure $5.85\pm 2.45$ net counts in the medium band (1 - 2 keV) and $4\pm 2$ counts in the hard band (2-7 keV). This hard photometry can be indicative of a significant level of obscuration, as no counts can be measured in the restframe 1.34 - 6.7 keV energy of the quasar SED.
In summary, we have presented evidence that supports that 1eRASS J202039.8-621525 is associated with the quasar emission, as (A) the Chandra sources ch2 and ch3 are marginally detected and do not overlap with the eRASS contours. (B) These sources are also not significantly detected in DES and CatWISE2020. (C) Among 151 potential DESI Legacy Survey candidates within 1’ of 1eRASS J202039.8-621525, the candidate with the highest relative probability p_i of being associated to the X-ray source is matched at the sub-arcsecond level with quasar coordinates. (D) No other Chandra source is detected to the southwest of the quasar, close to eROSITA events. While we cannot exclude the possibility of a transient phenomenon in a foreground galaxy, we conclude that it is likely that the bright eRASS1 source is associated with the quasar.
Table 4: Chandra ACIS-S photometry for three sources detected within 30” of the quasar.
ID
@@
$[\sigma]$
$F_{\mathrm{0.5-7.0keV}}$
[$\mathrm{erg\,s^{-1}\,cm^{-2}}$]
ch1 0.30 $9.9\pm 3.2$ 3.06 $5.83^{+3.70}_{-2.63}$
ch2 21.57 $4.7\pm 2.2$ 2.00 $2.86^{+2.98}_{-1.82}$
ch3 29.96 $6.6\pm 2.7$ 2.40 $3.80^{+3.20}_{-2.09}$
101010[1] Chandra source [2] The separation in arcsec between the optical position and the Chandra source. [3] The net counts in the band 0.5-7 keV [4] The srcflux-based significance of the detection based on source and background counts. [5] The flux in observed 0.5-7 keV.
@@ 4.3.2 Long-term X-ray variability
The eROSITA image presented in Fig. 11 displays two distinct lobes, which extend to the south and southwest of the coordinates of the quasar. One of these lobes coincides astrometrically with the quasar. In the Chandra image, three sources were detected with the wavdetect algorithm. One of these sources is matched at the sub-arcsecond level to the quasar’s coordinates. The two other sources are detected at a lower significance level to the north and the northwest of the quasar. Their location does not match the eRASS contours, and they are associated with DES sources within 1”. Assuming that the eRASS1 photometry is solely associated with the quasar, the eRASS and Chandra photometry presented respectively in Section 4.1 and 4.3.1, imply a decrease in the measured 0.5 - 7 keV flux of the source of about an order of magnitude from $F_{0.5-7\rm keV}=7.82^{+11.06}_{-4.47}\times 10^{-14}\mathrm{erg\,s^{-1}cm^{-2}}$ to $F_{0.5-7\rm keV}=5.83^{+3.70}_{-2.63}\times 10^{-15}\mathrm{erg\,s^{-1}cm^{-2}}$. We note that the flux decrease is only significant at the $\sim 2\sigma$ level due to the low number of net source counts in eRASS1. The two observations are separated by 3.5 years in the observed frame, corresponding to 6 months in the quasar restframe.
Fig. 12 shows the flux-decrease in the broad 0.5-7 keV band. The BXA flux posteriors for the eRASS1 and Chandra, their median and 15.9th and 84.1th percentiles are shown. We have further derived $L_{2-10\,\mathrm{keV}}$ and $\alpha_{OX}$ from the this Chandra flux posteriors. They are respectively marked in the left and right panels of Fig. 8.
@@ 4.3.3 Fast outflows
Here $f(v)$ denotes the continuum-normalized flux, $C$ is a constant equal to 1 for velocities $>2000\,\mathrm{km\,s^{-1}}$ with respect to the trough edge and 0 elsewhere. The measurement of the continuum is challenging due to the numerous BALs. We have approximated the continuum by using the principal component decomposition of $z\sim 3$ SDSS quasar continua presented by Pâris et al. (2011). The spectrum is strongly attenuated toward longer wavelengths, which is not accounted for by the PCA model. We fitted an additional power-law in the restframe spectral regions 1300-1350 $\AA$ and 1400-1600 $\AA$ to the spectrum normalized by the PCA-reprojected model. Finally, the spectrum of the source was normalised by the combined PCA and power-law models. We account for the flux uncertainties by bootstrapping from the spectrum RMS. From the normalized spectrum we measure $BI_{\mathrm{CIV}}$ = $3394\pm 197\,\mathrm{km\,s^{-1}}$ and $BI_{\mathrm{SiIV}}$ = $2935\pm 110\,\mathrm{km\,s^{-1}}$. This BAL strength confirms the classification of J202040-621509 as BAL quasar.
The large BALs are detached with respect to the quasar redshift, i.e. the troughs of all species are shifted shortwards of the expected corresponding quasar emission lines. We measure detached and terminal velocities from the start and the end of the absorption troughs. From the SiIV BAL we measure outflow velocities $v_{\mathrm{SiIV}}=0.002c-0.020c$ and $v_{\mathrm{CIV}}=0.003c-0.025c$. We note that the optical spectrum and the X-ray data were not obtained simultaneously. The LDSS3 observations took place closer to the eRASS1 observation ($\sim 1$ year in the observed frame) than the Chandra ACIS-S pointing ($\sim 2$ years), i.e. closer to the high-luminosity state discussed in Section 4.3.2.
@@ 4.4 Radio properties
J202040--621509 is detected at radio frequencies, at 0.887 and 1.37 GHz in VAST and at 0.943 GHz in the EMU Pilot survey (Norris et al. 2021). We retrieved the corresponding radio images from CASDA and we performed a single Gaussian fit using the task IMFIT within the CASA software to quantify the flux densities (see Table 6). The images are shown in Fig. 15 (top two rows) in Appendix C. Several VAST images are available, corresponding to different observing dates. We extracted the flux density from each of them. By assuming a single power-law for the continuum radio emission, we obtained a radio spectral index ($\alpha_{r}$) of -0.07$\pm$0.15 (see Fig. 14) that makes J202040--621509 FSRQ, similar to J020916--562650. The VAST measurements are consistent with each other, indicating that the source does not appear to be variable on time scales of a few months (observed frame). Then we computed the values of radio-loudness (R), by following the same procedure as for J020916--562650. We obtained a mean R of 10$\pm$4 which places the source at the boundary between objects classified as radio-quiet or radio-loud (Urry & Padovani 1995). This could indicate the presence of a relativistic jet (e.g., Sbarrato et al. 2021).
@@
4.5 Obscuring fast accretion disc winds at high redshift and other variability scenarios
@@
AGN shows substantial variability in flux and spectral shape at all frequencies (for a review, see Lira 2021). This variability arises from accretion physics (e.g. Nandra et al. 1997), broad line region kinematics (e.g. Cherepashchuk & Lyutyi 1973; Boksenberg et al. 1978; Blandford & McKee 1982; Peterson 1993) and jets (Albert et al. 2007, e.g.). In X-rays, AGN has been observed to be variable on timescales of hours to days (e.g. Winkler & White 1975; Marshall et al. 1981; McHardy et al. 2004; Ponti et al. 2012). Little redshift evolution in the X-ray variabity of AGN, as traced by the variance-frequency diagram, has been observed from $z=0$ to $z=3$ (Paolillo et al. 2023). Interpreted as reverberation effect, i.e. reprocessing of accretion disk photons in the electron corona (Haardt & Maraschi 1991), rapid X-ray variability can be used to study the structure of the innermost region of the AGN at a couple of gravitational radii away from the black hole (e.g. McHardy 2013). Coronal variability usually shows maximum amplitudes of a factor $\sim 2$ (e.g., Gibson & Brandt 2012). However, more extreme variability, unrelated to coronal fluctuations, has been observed on longer timescales (e.g. Timlin et al. 2020). For instance, (LaMassa et al. 2015) reported the first discovery of a changing look quasar in which the AGN flux dropped by a factor of 6 over timescales of years while not showing absorption signatures. As presented in Section 4.3.3, the UV spectrum of J202040-621509, in contrast, is that of a typical BAL quasar. The X-ray variability in quasars containing BALs in their restframe UV spectra is thought to arise from absorption effects. More precisely, variations in the column density of intervening, obscuring clouds are expected to arise from large-scale outflows (e.g. Brandt et al. 2000; Green et al. 2001; Gallagher et al. 2002, 2006; Gibson et al. 2009; Saez et al. 2012). An often invoked physical model for the observed extreme outflow velocities is that of line-driven winds launched from the accretion disk, as high-accretion rates drive the UV radiation pressure of the disk up. High-density regions, located at the inner edge of this flow can shield the coronal X-ray emission through photo-electric absorption (NH$>10^{22}\,\rm cm^{-2}$ Murray et al. 1995).
@@
A possible scenario is that the observed X-ray dimming of J202040-621509 over one order of magnitude in X-ray flux may partly be explained by such obscuring winds. While too few counts have been collected in the Chandra observation to extract and fit an X-ray spectrum, we only collected counts at energies $>6.7\,\rm keV$ in the quasar rest frame. This hard photometry can, for example, be indicative of heavy X-ray absorption, affecting lower-energetic photons.
@@
Extreme X-ray variability in high-redshift quasars has been reported previously (e.g. at $z=5.41$, Shemmer et al. 2005). Within the first gigayear, only two objects were known to display such a behaviour (Nanni et al. 2018; Vito et al. 2022). Nanni et al. (2018) reported a hardening of the spectrum and a flux decrease of factor 2.5 in the quasar SDSS J1030+0524 ($z=6.31$) between two XMM-Newton observations spaced by 14 years. They proposed intrinsic variation of accretion rate as a potential driver of the observed variability. Vito et al. (2022) presented the surprising non-detection in a 100ks XMM-Newton observations of the quasar CFHQS J164121+375520 ($z=6.047$), which had been detected 115 quasar-restframe days before with Chandra, implying a flux-dimming of factor $>7$. The authors discuss intervening obscuring material on sub-pc scales, such as disk-shape variability in super-Eddington sources. Vito et al. (2022) also list a possible model by Giustini & Proga (2019), in which rapidly accreting black holes with $\dot{m}\geq 0.25$, generate the favourable conditions for their optically-thick, thin disk emission to dominate the output of a cooled corona. These disks are expected to launch strong line-driven winds, imprinting BAL features in the spectrum of the quasars.
@@
While the latter model yields a promising scenario for the phenomenology observed in J202040-621509, we cannot exclude any of the above-mentioned sources of variability as the counting statistics of our X-ray observations are poor and we are missing simultaneous UV photometry to simultaneously track the corona and disk emission. A deeper, simultaneous X-ray and UV monitoring campaign will be required to shed light on this seemingly extreme X-ray dimming event.
@@ 5 Conclusions
In this work, we have presented a pilot survey designed to uncover new X-ray luminous quasars at $z>5.6$, combining eROSITA data with optical and IR photometry from DES, VHS and CatWISE2020. We have discovered 5 quasars out of 9 sources that were selected for spectroscopic confirmation with LDSS3-C. Two of these sources are securely detected in eRASS. These sources are also radio-detected. J020916-562650 is formally radio-loud, while J202040-621509 lies right at the threshold delimiting radio-loud from radio-quiet sources ($R=10$). The reionisation-era quasars uncovered by our pilot study shed light on powerful X-ray emission mechanisms in the early universe and give us access to demographic diagnostics for the $z>5.6$ quasar population.
J020916-562650 ($z=5.6$) shows X-ray and radio properties of a blazar, making it one of the most distant AGN of this type known to date (Belladitta et al. 2020; An et al. 2023; Caccianiga et al. 2024). In their independent analysis, Ighina et al. (2024) also come to the conclusion that this source is a blazar. Recently, the discovery of a blazar at $z=7$ was reported by Banados et al. (2024), setting a new redshift frontier for this type of source. From this detection and using simple geometric arguments, we infer a space density of similarly luminous, jetted quasars of $6.89^{15.86}_{-5.70}\,\mathrm{Gpc^{-3}}$. This value is consistent with the space density of all quasars, jetted or not, as measured by the most recent quasar luminosity functions at this redshift and brightness. Our measurement implies two potential scenarios: either all quasars discovered so far at $z\sim 6$ and M1450¡-25 are jetted (and most of these jets have not been detected yet) or a substantial fraction of jetted quasars in this redshift-magnitude regime have eluded quasar searches so far, perhaps due to obscuration. Severe systematics that can affect our measurement of quasar space density are discussed in this paper.
J202040-621509, ($z=5.7$) is an $R=10$ BAL quasar. Through four consecutive eRASS observations and an additional Chandra pointing, we provide substantial evidence for an extreme X-ray dimming event (a factor $\geq 13$ in X-ray flux). Potential dimming scenarios are nuclear obscuration or disk-dominated quasar SEDs in strongly outflowing quasars with rapidly accreting black holes. We conclude, however, that increased X-ray counting statistics and simultaneous UV monitoring will be required to draw clearer conclusions on this dimming event.